OpenStudy (anonymous):

Find the geometric mean of the pair of numbers: 81 and 4. Question 18 options: a) 38 b) 23 c) 28 d) 18

OpenStudy (amistre64):

take the square root of the product ...

OpenStudy (amistre64):

the geoMean of a and b is: sqrt(ab)

OpenStudy (lacypennelll):

@amistre64 is right you just need to take the square root and then poof you get the answer

OpenStudy (lacypennelll):

Ok if you really want the answer, look at other questions with different examples to figure it out.

OpenStudy (phi):

amistre told you how to do it. find the square root of the product ( product means multiply the two numbers)

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (phi):

can you multiply 81 and 4 ?

OpenStudy (anonymous):

ive asked what the two numbers are but they wont tell me are the 81 and 4?

OpenStudy (anonymous):

oh it is

OpenStudy (anonymous):

81 x 4 is 324

OpenStudy (phi):

yes, the sentence Find the geometric mean of the pair of numbers: 81 and 4. means use 81 and 4 now find the square root of 324 you should learn an important idea: factor the number into prime factors, and look for pairs of numbers. it is easiest if you start with 4*81 (don't multiply it out to get 324) instead, factor 4 into 2*2 and 81 into 3*3*3*3 so you have sqrt(2*2*3*3*3*3) pull out pairs of numbers

OpenStudy (anonymous):

ok

OpenStudy (phi):

sqrt(2*2*3*3*3*3) = sqrt(2*2) * sqrt(3*3) * sqrt(3*3) and sqrt(2*2)= 2 sqrt(3*3)= 3 sqrt(3*3)= 3

OpenStudy (anonymous):

im still lost a little

OpenStudy (phi):

do you agree 324 = 2*2*3*3*3*3 ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

do you know that sqrt( A * A) = A you can simplify two A's inside a square root by making them 1 A outside the square root

OpenStudy (phi):

that rule works for sqrt(2*2*3*3*3*3) notice you can write that as sqrt( (2*3*3) * (2*3*3) ) by writing the numbers as pairs you can simplify that to sqrt( (2*3*3) * (2*3*3) ) = 2*3*3

OpenStudy (phi):

or you can say: for every pair inside the square root, I can take that pair out, and put one of the pair outside the square root so : sqrt(2*2*3*3*3*3) = 2 sqrt(3*3*3*3) we took out the 2*2 from inside, and put a single 2 on the outside we can to that for the 3's also: 2 * 3* sqrt(3*3) we took out 1 pair of 3's we can take out the last pair of 3's 2*3*3*sqrt(1) = 2*3*3= 18

OpenStudy (anonymous):

ok

OpenStudy (phi):

or use a calculator on sqrt(324) (but it would be good to be able to do it without a calculator)

OpenStudy (anonymous):

the sqrt is 18

OpenStudy (phi):

yes, that means the geometric mean of 4 and 81 is 18

OpenStudy (anonymous):

ok Thanks :DD

OpenStudy (anonymous):

i have 5 more questions can you help me with those also?